OpenStudy (anonymous):

Express 45 = x as a logarithmic equation. @ranga

OpenStudy (anonymous):

log_45 (x) = 1

OpenStudy (anonymous):

Thats not an option

OpenStudy (anonymous):

give me the options

OpenStudy (ranga):

There are a few different ways of expressing it as a logarithmic equation. log(45) = log(x) is one and there are others: log(x/45) = 0 log(45/x) = 0 etc. So if you give the options we can help you select the right one.

OpenStudy (anonymous):

log45 = x log4x = 5 log54 = x log5x = 4

OpenStudy (ranga):

Or could you post a picture of the problem and the choices?